A pressure cooker contains 0.5m^3 water and water vapour mixture at 300 Celsius. Calculate the mass of each if their volumes are equal.
1
Expert's answer
2019-04-04T09:03:06-0400

From the steam table at "300 ^\circ C" (https://www.tlv.com/global/TI/calculator/steam-table-temperature.html):


"v_f = 0.00140422 \frac{m^3}{kg}"

"v_g = 0.0216631 \frac{m^3}{kg}"

(i) Calculate the mass of water:


"V_L = \frac{0.500 m^3}{2} = 0.250 m^3"


"v_f = \frac{V_L}{m_L}= 0.00140422 \frac{m^3}{kg}"


"\therefore m_L = \frac{V_L}{v_f} = \frac{0.250 m^3}{0.00140422 \frac{m^3}{kg}} = 178 kg"


(ii) Calculate the mass of water vapour mixture:


"V_v = \frac{0.500 m^3}{2} = 0.250 m^3"


"v_g = \frac {V_v}{m_v} = 0.0216631 \frac{m^3}{kg}"


"\therefore m_v = \frac{0.250 m^3}{0.0216631 \frac{m^3}{kg}}= 11.54 kg"


Answer: (i) "m_L = 178 kg"

(ii) "m_v = 11.54 kg"